i start walking at 11am and reach my destination by 1pm rowing against the stream.On coming back if start at 2pm, i reach rhe starting point by 3.What is the ratio of speed of the boat to that of stream
Answer by stanbon(75887)

i start walking at 11am and reach my destination by 1pm rowing against the stream.On coming back if start at 2pm, i reach rhe starting point by 3.What is the ratio of speed of the boat to that of stream
---
Upstream DATA::
time = 2 hrs ; distance = x miles ; rate = x/2 mph
Downstream DATA:
time = 1 hr ; distance = x miles ; rate = x mph
----
Equations:
b - c = x/2
b + c = x
----
2b = (3/2)x
Speed of boat = (3/4)x
----
2c = (x/2)
Speed of current = (1/4)x
---
Ans:: Ratio of speed of boat to speed of current is 3:1
-------------
Cheers,
Stan H.
----------